# On diffusion-based generative models and their error bounds: The log-concave case with full convergence estimates

### v4

@@ Abstract
We provide full theoretical guarantees for the convergence behaviour of diffusion-based generative models under the assumption of strongly log-concave data distributions while our approximating class of functions used for score estimation is made of Lipschitz continuous functions avoiding any Lipschitzness assumption on the score function. We demonstrate via a motivating example, sampling from a Gaussian distribution with unknown mean, the powerfulness of our approach. In this case, explicit estimates are provided for the associated optimization problem, i.e. score approximation, while these are combined with the corresponding sampling estimates. As a result, we obtain the best known upper bound estimates in terms of key quantities of interest, such as the dimension and rates of convergence, for the Wasserstein-2 distance between the data distribution (Gaussian with unknown mean) and our sampling algorithm.
Beyond the motivating example and in order to allow for the use of a diverse range of stochastic optimizers, we present our results using an $L^{2}$-accurate score estimation assumption, which crucially is formed under an expectation with respect to the stochastic optimizer and our novel auxiliary process that uses only known information. This approach yields the best known convergence rate for our sampling algorithm.
@@ 1 Introduction
Diffusion-based generative models, also known as score-based generative models (SGMs) (Song and Ermon, 2019; Song et al., 2021; Sohl-Dickstein et al., 2015; Ho et al., 2020), have become over the past few years one of the most popular approaches in generative modelling due to their empirical successes in data generation tasks. These models have achieved state-of-the-art results in image generation (Dhariwal and Nichol, 2021; Rombach et al., 2022), audio generation (Kong et al., 2020) and inverse problems (Chung et al., 2022; Song et al., 2022; Cardoso et al., 2024; Boys et al., 2023) outperforming other generative models like generative adversarial networks (GANs) (Goodfellow et al., 2014), variational autoencoders (VAEs) (Kingma and Welling, 2014), normalizing flows (Rezende and Mohamed, 2015) and energy-based methods (Zhao et al., 2017).
@@
The widespread applicability and success of SGMs have been accompanied by a growing interest in the theoretical understandings of these models, particularly in their convergence theory e.g. in (Bortoli, 2022; Block et al., 2020; Lee et al., 2022, 2023; Chen et al., 2023a, d; Li et al., 2024; Conforti et al., 2023; Benton et al., 2024; Oko et al., 2023; Pedrotti et al., 2023). At its core, this new generative modeling approach combines optimisation and sampling procedures – specifically, the approximation of the score and the creation of new samples, which make its theoretical analysis both an interesting and a rich challenge.
Some of the recent advances in the study of the theoretical properties of SGMs concentrate around the sampling procedure by assuming suitable control for the score estimation procedure. For instance, the analysis in (Lee et al., 2022; Chen et al., 2023d) assumes that the score estimate is $L^{2}$-accurate, meaning that the $L^{2}$ error between the score and its estimate is small, and provides estimates in total variation (TV) distance. Under the same assumption, the more recent contribution (Conforti et al., 2023) establishes non-asymptotic bounds in Kullback Leibler (KL) divergence by assuming finite relative Fisher of data distribution with respect to a Gaussian distribution.
@@
Our approach introduces a novel auxiliary process that relies solely on known information and uses the density of the solution of this process in the $L^{2}$-accurate score estimation. To further highlight the powerfulness of our approach, we present a motivating example on the case of sampling from a Gaussian distribution with unknown mean. Full theoretical estimates for the convergence properties of the SGM are provided in Wasserstein-2 distance, while our choice of stochastic optimizer for the score approximation is a simple Langevin-based algorithm. Our estimates are explicit and contain the best known optimal dependencies in terms of dimension and rate of convergence (Theorem 3.1). To the best of the authors’ knowledge, these are the first such explicit results with transparent dependence on the parameters involved in the sampling and optimization combined procedures of the diffusion models. By connecting the diffusion models with the theoretical guarantees of machine learning optimizers via standard stochastic calculus tools, the results in Theorem 3.1, together with the bounds achieved in the more general setting in Theorem 3.10, provide the theoretical justification for the empirical success of the diffusion models.
We close this introductory section by highlighting some other, alternative approaches which were recently developed. One may consult (Yang and Wibisono, 2023) for an approach based on the assumption that the score estimation error has a sub-Gaussian tail. This is a stronger assumption than $L^{2}$-accurate. In (Chen et al., 2023b), non-asymptotic bounds in TV and Wassertein distance of order 2 are derived when the data distribution is supported on a low-dimensional linear subspace. Finally, convergence guarantees in TV were developed in (Chen et al., 2023c) for the probability flow ordinary differential equation (ODE) implementation of SGMs under the $L^{2}$-accurate score estimate assumption.
@@
$W_{2}(\mu,\nu)=\left(\inf_{\zeta\in\mathscr{C}(\mu,\nu)}\int_{\mathbb{R}^{M}}\int_{\mathbb{R}^{M}}|x-y|^{2}\ \text{d}\zeta(x,y)\right)^{\frac{1}{2}}.$
@@ 2 Technical Background
In this section, we provide a brief summary behind the construction of score-based generative models (SGMs) based on diffusion introduced in (Song et al., 2021). The fundamental concept behind SGMs centers around the use of an ergodic (forward) diffusion process to diffuse the unknown data distribution $\pi_{\mathsf{D}}\in\mathscr{P}(\mathbb{R}^{M})$ to a known prior distribution and then learn a backward process to transform the prior to the target distribution $\pi_{\mathsf{D}}$ by estimating the score function of the forward process. In our analysis, we focus on the forward process $(X_{t})_{t\in[0,T]}$ given by an Ornstein-Uhlenbeck (OU) process
@@
where $(B_{t})_{t\in[0,T]}$ is an $M$-dimensional Brownian motion and we assume that $\mathbb{E}[|X_{0}|^{2}]<\infty$. The process (1) is chosen to match the forward process in the original paper (Song et al., 2021), which is also referred to as Variance Preserving Stochastic Differential Equation. The noising process (1) can also be represented as follows
@@
$\text{d}Y_{t}=(Y_{t}+2\nabla\log p_{T-t}(Y_{t}))\ \text{d}t+\sqrt{2}\ \text{d}\bar{B}_{t},\quad Y_{0}\sim\mathscr{L}(X_{T}),$ (3)
where $\{p_{t}\}_{t\in[0,T]}$ is the family of densities of $\{\mathscr{L}(X_{t})\}_{t\in(0,T]}$ with respect to the Lebesgue measure, $\bar{B}_{t}$ is an another Brownian motion independent of $B_{t}$ in (1) defined on $(\Omega,\mathscr{F},\mathbb{P})$. However, the sampling is done in practice from the invariant distribution of the forward process, which, in this case, is a standard Gaussian distribution. Therefore, the backward process (3) becomes
$\text{d}\widetilde{Y}_{t}=(\widetilde{Y}_{t}+2\ \nabla\log p_{T-t}(\widetilde{Y}_{t}))\ \text{d}t+\sqrt{2}\ \text{d}\bar{B}_{t},\quad\widetilde{Y}_{0}\sim\pi_{\infty}=\mathscr{N}(0,I_{M}).$ (4)
@@
To account for numerical instability issues for training and sampling at $t=0$ as observed in practice in (Song et al., 2021, Appendix C) and for the possibility that the integral of the score function in (5) may diverge when $t=0$ (see Appendix A), a discretised version of the score-matching optimization problem is usually considered
@@
where $\epsilon>0$ and $\kappa:[0,T]\rightarrow\mathbb{R}_{>0}$ is a weighting function. The score-matching objective $U$ in (6) can be rewritten via denoising score matching (Vincent, 2011) as follows
@@
where the expectation is over $\tau\sim\text{Uniform}([\epsilon,T])$, $X_{0}\sim\pi_{\mathsf{D}}$ and $Z\sim\mathscr{N}(0,I_{M})$, and where $C\in\mathbb{R}$ is a constant independent of $\theta$ (see Appendix A for the derivation with the OU representation (26)). The stochastic gradient $H:\mathbb{R}^{d}\times\mathbb{R}^{m}\to\mathbb{R}^{d}$ of (6) deduced using (7) is given by
@@
$\text{d}Y_{t}^{\text{aux}}=(Y_{t}^{\text{aux}}+2\ s(T-t,\hat{\theta},Y_{t}^{\text{aux}}))\ \text{d}t+\sqrt{2}\ \text{d}\bar{B}_{t},\quad Y_{0}^{\text{aux}}\sim\pi_{\infty}=\mathscr{N}(0,I_{M}).$ (9)
The process (9) will play an important role in the derivation of the nonasymptotic estimates in Wasserstein distance of order two between the target data distribution and the generative distribution of the diffusion model. Indeed, it connects the backward process (4) and the numerical scheme (11), which facilitates the analysis of the convergence of the diffusion model (see Appendix B.2 and Appendix C.2 for more details). For this reason, we introduce a sequence of stepsizes $\{\gamma_{k}\}_{k\in\{0,\dots,K\}}$ such that $\sum_{k=0}^{K}\gamma_{k}=T$. For any $k\in\{0,\dots,K\}$, let $t_{k+1}=\sum_{j=0}^{k}\gamma_{j}$ with $t_{0}=0$ and $t_{K+1}=T$. Let $\gamma_{k}=\gamma\in(0,1)$ for each $k=0,\dots,K$. The discrete process $(Y_{k}^{\text{EM}})_{k\in\{0,\dots,K+1\}}$ of the Euler–Maruyama approximation of (9) is given, for any $k\in\{0,\dots,K\}$, as follows
$Y_{k+1}^{\text{EM}}=Y_{k}^{\text{EM}}+\gamma(Y_{k}^{\text{EM}}+2\ s(T-k,\hat{\theta},Y_{k}^{\text{EM}}))+\sqrt{2\gamma}\ \bar{Z}_{k+1},\quad Y_{0}^{\text{EM}}\sim\pi_{\infty}=\mathscr{N}(0,I_{M}),$ (10)
where $\{\bar{Z}_{k}\}_{k\in\{0,\dots,K+1\}}$ is a sequence of independent $M$-dimensional Gaussian random variables with zero mean and identity covariance matrix. We emphasize that the approximation (10) is the one chosen in the original paper (Song et al., 2021). Finally, the continuous-time interpolation of (10), for $t\in[0,T]$, is given by
$\mathrm{d}\widehat{Y}_{t}^{\text{EM}}=(\widehat{Y}_{\lfloor t/\gamma\rfloor\gamma}^{\text{EM}}+2\ s(T-\lfloor t/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}_{\lfloor t/\gamma\rfloor\gamma}^{\text{EM}}))\ \text{d}t+\sqrt{2}\ \text{d}\bar{B}_{t},\qquad\widehat{Y}_{0}^{\text{EM}}\sim\pi_{\infty}=\mathscr{N}(0,I_{M}),$ (11)
where $\mathscr{L}(\widehat{Y}_{k}^{\text{EM}})=\mathscr{L}(Y^{\text{EM}}_{k})$ at grid points for each $k\in\{0,\dots,K+1\}$.
@@ 3.1 A Motivating Example: Full Estimates for Multivariate Gaussian Initial Data with Unknown Mean
In this section, we consider the case where the data distribution follows a multivariate normal distribution with unknown mean and identity covariance, i.e., $X_{0}\sim\pi_{\mathsf{D}}=\mathscr{N}(\mu,I_{d})$ for some unknown $\mu\in\mathbb{R}^{d}$ with $M=d$. We show that, by using diffusion models, we are able to generate new data from an approximate distribution that is close to $\pi_{\mathsf{D}}$. More precisely, we provide a non-asymptotic convergence bound with explicit constants in Wasserstein-2 distance between the law of the diffusion model and $\pi_{\mathsf{D}}$, which can be made arbitrarily small by appropriately choosing key parameters on the upper bound. In this example, the estimation of the score function reduces to the estimation of the unknown mean by the methods of convex optimization with optimal dependence of the dimension combined with the most efficient sampling method for high-dimensional Gaussian data.
In this setting by using (1), we can derive the score function given by
@@
To obtain an approximated score, i.e., to obtain an optimal value of $\theta$ in (13), we opt for a popular class of algorithms called stochastic gradient Langevin dynamics (SGLD) to solve the optimisation problem (6). In addition, we choose the weighting function $\kappa(t)=\sigma_{t}^{2}$ as in (Song and Ermon, 2019) and we set $\epsilon=0$ in (6). Using the approximating function (13) in (8), we can obtain the following expression for the stochastic gradient
@@
is a sequence of i.i.d. random variables generated as follows. For each $n\in\mathbb{N}_{0}$, we sample $\tau_{n}$ from $\text{Uniform}([0,T])$ such that $\mathscr{L}(\tau_{n})=\mathscr{L}(\tau)$, sample $X_{0,n}$ from $\pi_{\text{D}}=\mathscr{N}(\mu,\text{I}_{d})$ such that $\mathscr{L}(X_{0,n})=\mathscr{L}(X_{0})$, and sample $Z_{n}$ from $\mathscr{N}(0,I_{d})$ such that $\mathscr{L}(Z_{n})=\mathscr{L}(Z)$. In addition, we consider the case where $\theta_{0}$, $(\xi_{n})_{n\in\mathbb{N}_{0}}$, and $(\mathbf{X}_{n})_{n\in\mathbb{N}_{0}}$ in (15) are independent and we have $\mathbb{E}[H(\theta,\mathbf{X}_{n+1})]=\nabla U(\theta)$.
@@
Table 1: Explicit expressions for the constants in Theorem 3.1.
Theorem 3.1 states the non-asymptotic (upper) bounds between the generative distribution of the diffusion model $\mathscr{L}(\widehat{Y}_{K+1}^{\text{EM}})$ and the data distribution $\pi_{\mathsf{D}}$. An overview of the proof can be found in Appendix B.2.
Theorem 3.1.
@@
$\begin{split}&W_{2}(\mathscr{L}(\widehat{Y}_{K+1}^{\text{EM}}),\pi_{\mathsf{D}})\\ &\leq\sqrt{2}e^{-2T}(\sqrt{\mathbb{E}\left[|X_{0}|^{2}\right]}+\sqrt{d})\\ &\quad+(\sqrt{4/3}+2\sqrt{33})(e^{-n\lambda\mathbb{E}[\sigma_{\tau}^{2}m_{\tau}^{2}]}\sqrt{\mathbb{E}[|\theta_{0}-\theta^{*}|^{2}]}+\sqrt{dC_{\mathsf{SGLD},1}/\beta}+\sqrt{\lambda C_{\mathsf{SGLD},2}})\\ &\quad+\gamma(\sqrt{18d}+\sqrt{132|\theta^{*}|^{2}}),\end{split}$ (18)
where $C_{\mathsf{SGLD},1}$ and $C_{\mathsf{SGLD},2}$ are given explicitly in Table 1. In addition, the result in (18) implies that for any $\delta>0$, if we choose $T>T_{\delta}$, $\beta\geq\beta_{\delta}$, $0<\lambda\leq\lambda_{\delta}$, $n\geq n_{\delta}$, and $0<\gamma<\gamma_{\delta}$, then
$W_{2}(\mathscr{L}(\widehat{Y}_{K+1}^{\text{EM}}),\pi_{\mathsf{D}})<\leavevmode\nobreak\ \delta,$
@@
Remark 3.2.
The result in Theorem 3.1 achieves the optimal rate of convergence of order one for Euler or Milstein schemes of SDEs with constant diffusion coefficients. Furthermore, one notes that the dependence of the dimension on the upper bound in (18) is in the form of $\sqrt{d}$. To the best of the authors’ knowledge, the result in Theorem 3.1 is the first convergence bound where the parameters involved in the sampling and optimization steps of the diffusion models appear explicitly. In the optimization procedure, we use SGLD algorithm (15) to solve the problem (6). Since the stochastic gradient $H$ in (14) is strongly convex by Proposition B.1, it has been proved, for instance in (Barkhagen et al., 2021), that, for large enough $\beta>0$, the output of SGLD is an almost minimizer of (6) when $n$ is large. Thus, in the diffusion model, we set $\hat{\theta}=\theta_{n}^{\lambda}$ indicating $\nabla\log p_{t}(x)\approx s(t,\hat{\theta},x)=s(t,\theta_{n}^{\lambda},x)$ for large values of $n$ and for all $t$ and $x$. Crucially, this allows us to use the established convergence results for SGLD to deduce a sampling upper bound for $W_{2}(\mathscr{L}(\widehat{Y}_{K+1}^{\text{EM}}),\pi_{\mathsf{D}})$ with explicit constants in (18). Consequently, this bound can be controlled by any given precision level $\delta>0$ by appropriately choosing $T,\beta,\lambda,n$ and $\gamma$.

### v5

@@ Abstract
We provide full theoretical guarantees for the convergence behaviour of diffusion-based generative models under the assumption of strongly log-concave data distributions while our approximating class of functions used for score estimation is made of Lipschitz continuous functions avoiding any Lipschitzness assumption on the score function. We demonstrate via a motivating example, sampling from a Gaussian distribution with unknown mean, the powerfulness of our approach. In this case, explicit estimates are provided for the associated optimization problem, i.e. score approximation, while these are combined with the corresponding sampling estimates. As a result, we obtain the best known upper bound estimates in terms of key quantities of interest, such as the dimension and rates of convergence, for the Wasserstein-2 distance between the data distribution (Gaussian with unknown mean) and our sampling algorithm. Beyond the motivating example and in order to allow for the use of a diverse range of stochastic optimizers, we present our results using an $L^{2}$-accurate score estimation assumption, which crucially is formed under an expectation with respect to the stochastic optimizer and our novel auxiliary process that uses only known information. This approach yields the best known convergence rate for our sampling algorithm.
@@ 1 Introduction
Diffusion-based generative models, also known as score-based generative models (SGMs) (Song & Ermon, 2019; Song et al., 2021; Sohl-Dickstein et al., 2015; Ho et al., 2020), have become over the past few years one of the most popular approaches in generative modelling due to their empirical successes in data generation tasks. These models have achieved state-of-the-art results in image generation (Dhariwal & Nichol, 2021; Rombach et al., 2022), audio generation (Kong et al., 2020) and inverse problems (Chung et al., 2022; Song et al., 2022; Cardoso et al., 2024; Boys et al., 2024) outperforming other generative models like generative adversarial networks (GANs) (Goodfellow et al., 2014), variational autoencoders (VAEs) (Kingma & Welling, 2014), normalizing flows (Rezende & Mohamed, 2015) and energy-based methods (Zhao et al., 2017).
@@
The widespread applicability and success of SGMs have been accompanied by a growing interest in the theoretical understandings of these models, particularly in their convergence theory e.g. in Block et al. (2020); De Bortoli et al. (2021); Bortoli (2022); Lee et al. (2022); Yang & Wibisono (2022); Kwon et al. (2022); Liu et al. (2022); Oko et al. (2023); Lee et al. (2023); Chen et al. (2023a; d); Li et al. (2024); Pedrotti et al. (2024); Conforti et al. (2025); Benton et al. (2024), with further works appearing after the first version of our preprint e.g. in Tang & Zhao (2024); Strasman et al. (2025); Mimikos-Stamatopoulos et al. (2024). At its core, this new generative modeling approach combines optimisation and sampling procedures – specifically, the approximation of the score and the creation of new samples, which make its theoretical analysis both an interesting and a rich challenge.
Some of the recent advances in the study of the theoretical properties of SGMs concentrate around the sampling procedure by assuming suitable control for the score estimation procedure. For instance, the analysis in Lee et al. (2022); Chen et al. (2023d) assumes that the score estimate is $L^{2}$-accurate, meaning that the $L^{2}$ error between the score and its estimate is small, and provides estimates in total variation (TV) distance. Under the same assumption, the more recent contribution in Conforti et al. (2025) establishes non-asymptotic bounds in Kullback Leibler (KL) divergence by assuming finite relative Fisher of data distribution with respect to a Gaussian distribution.
@@
Our approach introduces a novel auxiliary process that relies solely on known information and uses the density of the solution of this process in the $L^{2}$-accurate score estimation. To further highlight the powerfulness of our approach, we present a motivating example on the case of sampling from a Gaussian distribution with unknown mean. Full theoretical estimates for the convergence properties of the SGM are provided in Wasserstein-2 distance, while our choice of stochastic optimizer for the score approximation is a simple Langevin-based algorithm. Our estimates are explicit and contain the best known optimal dependencies in terms of dimension and rate of convergence (Theorem 1). To the best of the authors’ knowledge, these are the first such explicit results with transparent dependence on the parameters involved in the sampling and optimization combined procedures of the diffusion models. By connecting the diffusion models with the theoretical guarantees of machine learning optimizers via standard stochastic calculus tools, the results in Theorem 1, together with the bounds achieved in the more general setting in Theorem 10, provide the theoretical justification for the empirical success of the diffusion models.
We close this introductory section by highlighting some other, alternative approaches which were recently developed. One may consult Yang & Wibisono (2023) for an approach based on the assumption that the score estimation error has a sub-Gaussian tail. This is a stronger assumption than $L^{2}$-accurate. In Chen et al. (2023b), non-asymptotic bounds in TV and Wassertein distance of order 2 are derived when the data distribution is supported on a low-dimensional linear subspace. Finally, convergence guarantees in TV were developed in Chen et al. (2023c) for the probability flow ordinary differential equation (ODE) implementation of SGMs under the $L^{2}$-accurate score estimate assumption.
@@
$W_{2}(\mu,\nu)=\left(\inf_{\zeta\in\mathcal{C}(\mu,\nu)}\int_{\mathbb{R}^{M}}\int_{\mathbb{R}^{M}}|x-y|^{2}\ \text{d}\zeta(x,y)\right)^{\frac{1}{2}}.$
@@ 2 Technical Background
In this section, we provide a brief summary behind the construction of score-based generative models (SGMs) based on diffusion introduced in Song et al. (2021). The fundamental concept behind SGMs centers around the use of an ergodic (forward) diffusion process to diffuse the unknown data distribution $\pi_{\mathsf{D}}\in\mathcal{P}(\mathbb{R}^{M})$ to a known prior distribution and then learn a backward process to transform the prior to the target distribution $\pi_{\mathsf{D}}$ by estimating the score function of the forward process. In our analysis, we focus on the forward process $(X_{t})_{t\in[0,T]}$ given by an Ornstein-Uhlenbeck (OU) process
@@
where $(B_{t})_{t\in[0,T]}$ is an $M$-dimensional Brownian motion and we assume that $\mathbb{E}[|X_{0}|^{2}]<\infty$. The process 1 is chosen to match the forward process in the original paper (Song et al., 2021), which is also referred to as Variance Preserving Stochastic Differential Equation. The noising process 1 can also be represented as follows
@@
$\text{d}Y_{t}=(Y_{t}+2\nabla\log p_{T-t}(Y_{t}))\ \text{d}t+\sqrt{2}\ \text{d}\bar{B}_{t},\quad Y_{0}\sim\mathcal{L}(X_{T}),$ (3)
where $\{p_{t}\}_{t\in[0,T]}$ is the family of densities of $\{\mathcal{L}(X_{t})\}_{t\in(0,T]}$ with respect to the Lebesgue measure, $\bar{B}_{t}$ is an another Brownian motion independent of $B_{t}$ in 1 defined on $(\Omega,\mathcal{F},\mathbb{P})$. However, the sampling is done in practice from the invariant distribution of the forward process, which, in this case, is a standard Gaussian distribution. Therefore, the backward process 3 becomes
$\text{d}\widetilde{Y}_{t}=(\widetilde{Y}_{t}+2\ \nabla\log p_{T-t}(\widetilde{Y}_{t}))\ \text{d}t+\sqrt{2}\ \text{d}\bar{B}_{t},\quad\widetilde{Y}_{0}\sim\pi_{\infty}=\mathcal{N}(0,I_{M}).$ (4)
@@
To account for numerical instability issues for training and sampling at $t=0$ as observed in practice in Song et al. (2021, Appendix C) and for the possibility that the integral of the score function in 5 may diverge when $t=0$ (see Appendix A), a discretised version of the score-matching optimization problem is usually considered
@@
where $\epsilon>0$ and $\kappa:[0,T]\rightarrow\mathbb{R}_{>0}$ is a weighting function. The score-matching objective $U$ in 6 can be rewritten via denoising score matching (Vincent, 2011) as follows
@@
where the expectation is over $\tau\sim\text{Uniform}([\epsilon,T])$, $X_{0}\sim\pi_{\mathsf{D}}$ and $Z\sim\mathcal{N}(0,I_{M})$, and where $C\in\mathbb{R}$ is a constant independent of $\theta$ (see Appendix A for the derivation with the OU representation 30). The stochastic gradient $H:\mathbb{R}^{d}\times\mathbb{R}^{m}\to\mathbb{R}^{d}$ of 6 deduced using 7 is given by
@@
$\text{d}Y_{t}^{\text{aux}}=(Y_{t}^{\text{aux}}+2\ s(T-t,\hat{\theta},Y_{t}^{\text{aux}}))\ \text{d}t+\sqrt{2}\ \text{d}\bar{B}_{t},\quad Y_{0}^{\text{aux}}\sim\pi_{\infty}=\mathcal{N}(0,I_{M}).$ (9)
The process 9 will play an important role in the derivation of the nonasymptotic estimates in Wasserstein distance of order two between the target data distribution and the generative distribution of the diffusion model. Indeed, it connects the backward process 4 and the numerical scheme 11, which facilitates the analysis of the convergence of the diffusion model (see Appendix C.2 and Appendix D.2 for more details). For this reason, we introduce a sequence of stepsizes $\{\gamma_{k}\}_{k\in\{0,\dots,K\}}$ such that $\sum_{k=0}^{K}\gamma_{k}=T$. For any $k\in\{0,\dots,K\}$, let $t_{k+1}=\sum_{j=0}^{k}\gamma_{j}$ with $t_{0}=0$ and $t_{K+1}=T$. Let $\gamma_{k}=\gamma\in(0,1)$ for each $k=0,\dots,K$. The discrete process $(Y_{k}^{\text{EM}})_{k\in\{0,\dots,K+1\}}$ of the Euler–Maruyama approximation of 9 is given, for any $k\in\{0,\dots,K\}$, as follows
$Y_{k+1}^{\text{EM}}=Y_{k}^{\text{EM}}+\gamma(Y_{k}^{\text{EM}}+2\ s(T-t_{k},\hat{\theta},Y_{k}^{\text{EM}}))+\sqrt{2\gamma}\ \bar{Z}_{k+1},\quad Y_{0}^{\text{EM}}\sim\pi_{\infty}=\mathcal{N}(0,I_{M}),$ (10)
where $\{\bar{Z}_{k}\}_{k\in\{0,\dots,K+1\}}$ is a sequence of independent $M$-dimensional Gaussian random variables with zero mean and identity covariance matrix. We emphasize that the approximation 10 is the one chosen in the original paper Song et al. (2021). Finally, the continuous-time interpolation of 10, for $t\in[0,T]$, is given by
$\mathrm{d}\widehat{Y}_{t}^{\text{EM}}=(\widehat{Y}_{\lfloor t/\gamma\rfloor\gamma}^{\text{EM}}+2\ s(T-\lfloor t/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}_{\lfloor t/\gamma\rfloor\gamma}^{\text{EM}}))\ \text{d}t+\sqrt{2}\ \text{d}\bar{B}_{t},\qquad\widehat{Y}_{0}^{\text{EM}}\sim\pi_{\infty}=\mathcal{N}(0,I_{M}),$ (11)
where $\mathcal{L}(\widehat{Y}_{k}^{\text{EM}})=\mathcal{L}(Y^{\text{EM}}_{k})$ at grid points for each $k\in\{0,\dots,K+1\}$.
@@ 3.1 A Motivating Example: Full Estimates for Multivariate Gaussian Initial Data with Unknown Mean
In this section, we consider the case where the data distribution follows a multivariate normal distribution with unknown mean and identity covariance, i.e., $X_{0}\sim\pi_{\mathsf{D}}=\mathcal{N}(\mu,I_{d})$ for some unknown $\mu\in\mathbb{R}^{d}$ with $M=d$. We show that, by using diffusion models, we are able to generate new data from an approximate distribution that is close to $\pi_{\mathsf{D}}$. More precisely, we provide a non-asymptotic convergence bound with explicit constants in Wasserstein-2 distance between the law of the diffusion model and $\pi_{\mathsf{D}}$, which can be made arbitrarily small by appropriately choosing key parameters on the upper bound. In this example, the estimation of the score function reduces to the estimation of the unknown mean by the methods of convex optimization with optimal dependence of the dimension combined with the most efficient sampling method for high-dimensional Gaussian data.
In this setting by using 1, we can derive the score function given by
@@
To obtain an approximated score, i.e., to obtain an optimal value of $\theta$ in 13, we opt for a popular class of algorithms called stochastic gradient Langevin dynamics (SGLD) to solve the optimisation problem 6. In addition, we choose the weighting function $\kappa(t)=\sigma_{t}^{2}$ as in Song & Ermon (2019) and we set $\epsilon=0$ in 6. Using the approximating function 13 in 8, we can obtain the following expression for the stochastic gradient
@@
is a sequence of i.i.d. random variables generated as follows. For each $n\in\mathbb{N}_{0}$, we sample $\tau_{n}$ from $\text{Uniform}([0,T])$ such that $\mathcal{L}(\tau_{n})=\mathcal{L}(\tau)$, sample $X_{0,n}$ from $\pi_{\text{D}}=\mathcal{N}(\mu,\text{I}_{d})$ such that $\mathcal{L}(X_{0,n})=\mathcal{L}(X_{0})$, and sample $Z_{n}$ from $\mathcal{N}(0,I_{d})$ such that $\mathcal{L}(Z_{n})=\mathcal{L}(Z)$. In addition, we consider the case where $\theta_{0}$, $(\xi_{n})_{n\in\mathbb{N}_{0}}$, and $(\mathbf{X}_{n})_{n\in\mathbb{N}_{0}}$ in 15 are independent and we have $\mathbb{E}[H(\theta,\mathbf{X}_{n+1})]=\nabla U(\theta)$.
@@
Theorem 1 states the non-asymptotic (upper) bounds between the generative distribution of the diffusion model $\mathcal{L}(\widehat{Y}_{K+1}^{\text{EM}})$ and the data distribution $\pi_{\mathsf{D}}$. An overview of the proof can be found in Appendix C.2.
Theorem 1.
@@
$\begin{split}&W_{2}(\mathcal{L}(\widehat{Y}_{K+1}^{\text{EM}}),\pi_{\mathsf{D}})\\ &\leq\sqrt{2}e^{-2T}(\sqrt{\mathbb{E}\left[|X_{0}|^{2}\right]}+\sqrt{d})\\ &\quad+(\sqrt{4/3}+2\sqrt{33})(e^{-n\lambda\mathbb{E}[\sigma_{\tau}^{2}m_{\tau}^{2}]}\sqrt{\mathbb{E}[|\theta_{0}-\theta^{*}|^{2}]}+\sqrt{dC_{\mathsf{SGLD},1}/\beta}+\sqrt{\lambda C_{\mathsf{SGLD},2}})\\ &\quad+\gamma(\sqrt{18d}+\sqrt{132|\theta^{*}|^{2}}),\end{split}$ (18)
where $C_{\mathsf{SGLD},1}$ and $C_{\mathsf{SGLD},2}$ are given explicitly in Table 1. In addition, the result in 18 implies that for any $\delta>0$, if we choose $T>T_{\delta}$, $\beta\geq\beta_{\delta}$, $0<\lambda\leq\lambda_{\delta}$, $n\geq n_{\delta}$, and $0<\gamma<\gamma_{\delta}$, then
$W_{2}(\mathcal{L}(\widehat{Y}_{K+1}^{\text{EM}}),\pi_{\mathsf{D}})<~{}\delta,$
@@
Remark 2.
The result in Theorem 1 achieves the optimal rate of convergence of order one for Euler or Milstein schemes of SDEs with constant diffusion coefficients. Furthermore, one notes that the dependence of the dimension on the upper bound in 18 is in the form of $\sqrt{d}$. To the best of the authors’ knowledge, the result in Theorem 1 is the first convergence bound where the parameters involved in the sampling and optimization steps of the diffusion models appear explicitly. In the optimization procedure, we use SGLD algorithm 15 to solve the problem 6. Since the stochastic gradient $H$ in 14 is strongly convex by Proposition 13, it has been proved, for instance in Barkhagen et al. (2021), that, for large enough $\beta>0$, the output of SGLD is an almost minimizer of 6 when $n$ is large. Thus, in the diffusion model, we set $\hat{\theta}=\theta_{n}^{\lambda}$ indicating $\nabla\log p_{t}(x)\approx s(t,\hat{\theta},x)=s(t,\theta_{n}^{\lambda},x)$ for large values of $n$ and for all $t$ and $x$. Crucially, this allows us to use the established convergence results for SGLD to deduce a sampling upper bound for $W_{2}(\mathcal{L}(\widehat{Y}_{K+1}^{\text{EM}}),\pi_{\mathsf{D}})$ with explicit constants in 18. Consequently, this bound can be controlled by any given precision level $\delta>0$ by appropriately choosing $T,\beta,\lambda,n$ and $\gamma$.